No need for Explanation

The CFO of a major corporation is trying to decide on what project to pursue using different investment criteria: Net Present Value (NPV), Payback Period, Discounted Payback Period, Internal Rate of Return (IRR), and the Profitability Index (PI). The CFO has four projects to choose between: Project W (which is a strip mine - see problem 6), Project X, Project Y, and Project Z. Additional information about each project is summarized below. You can use the workspace provided to help answer questions 6-11: Project W Costs $900M today (T-0) The appropriate discount rate is 6.0% Project w generates six after tax cash flows beginning one year from now ( 1); the value of each cash flow is in the table below YEAR 0 10 500 S 200 ProjectX Costs $800M today (1-0) The appropriate discount rate is 6.0% Poect X generates six equal $240M after tax cash flows begihning one year from now (T-1) YEAR Project X CF1 CF2 CF3 CF4 CF5 CF6 Project Y Costs $1000M today (T-0) The appropriate discount rate is 8.0% Project Y generates infinite after tax cash flows; the first cash flow is $60M one year from now (T=1); cash flows thereafter will grow at 4.0% per year forever YEAR Project Y CF1 CF2 CF3 CF4 CF5 CF6 CF8 Project Z Costs $1200M today (T-0) The appropriate discount rate is 7.0% Project Z generates 30 equal $120M after tax cash flows: the first cash flow is one year from now (T-1) YEAR 0 30 5 CF5 Project Z CF1 CF2 CF3 CF4 CF6 CF8 CF306. Project W is a strip mine which requires you to clean up environmental damage when retiring the mine; this is why cash flows in year 5 and year 6 are negative. The NPV of Project W is closest to: a. $442 b. $600 c. $640 d. $1440 e. $1500 7. The CFO decides to pursue only one project, but the project is required to have an NPV greater than or equal to $500M. If more than one project satisfies this criteria, the CFO will make the decision to pick the project with the highest Pl. Which project would you choose: a. W b. X C. d. Z e. None of the projects 8. For this problem only, assume the CFO only has $2000 to invest today, the CFO cannot borrow any additional funds, and any remaining funds (not invested in the projects) earns 0.0% interest. The CFO wants to maximize NPV by pursuing as many projects possible. Which project (s) would the CFO choose: a. W, X, Y, Z d. X & Z e. Y & Z 9. The IRR of Project X is closest to: a. 10.0% b. 12.5% 15.0% 20.0% 25.0% d. e. 10.For this problem only, assume the CFO only has $2000 to invest today, the CFO cannot borrow any additional funds, and any remaining funds (not invested in the projects) earns 0.0% interest. The CFO is considering picking the project with the highest IRR. You are hired as a consultant; how would you advise the CFO to proceed? a. Pick Project W because it has the highest IRR b. Pick Project X because it has the highest IRR c. Pick Project Y because it has the highest IRR d. Dont use IRR to rank mutually exclusive projects, use a different criteria to pick the best project e. Dont pick Project W because it has unconventional cash flows; instead, Pick Project X because it has the second highest IRR 11. For Project X, the discounted payback period is about longer than the payback period a. 3 months b. 4 months c. 6 months d. 7 months e. 10 months
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins ー E ゴWrap Text B 1 u. ,_a. ars-函Merge & Center, $, % , 弼,8 C Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard D144 PV FACTOR AT 6% PV OF PROJECT W YEAR PROJECT X NPV = PV OF ALL ACSH INFLOWS-INVESTMENT NPV (240 X PVI FA@6%, 6 YEARS)-800 NPV(240 X4.9173) - 800 0.9434 283.02 0.8900 445.00 0.8396671.70 0.7921 158.42 0.747374.73 0.7050 -140.99 442.41 380.15 500 PROJECTY PERPETUAL CASHFLOWS WITH GROWTH RATE PV = PV OF ALL ACSH IN FLOWS-INVESTMENT NPVCF/(r -g) -INVESTMENT NPV60/(0.08-0.04) 1000 500.00 PROJECTz NPV-PV OF ALL ACSH INFLOWS INVESTMENT NPV- (120XPVIFA@7%, З0 YEARS)-1200 NPV= (120 X 12.4090)-1200 289.08 IRRmpv PBP erences: E147 14 external rate MIRR NpV IRR REPLACEMENT S i HPR GMAM EAC MACRS LEASE CASH BUDGET i BOND EPS EBIT REPLACEMENT ( vacc bond sens tvt i ALE 02391 0074794 3518402 7574 PCA 9275 988777 |- 58212 0123456 45 46 47 48 49 - 23 56789 1 2 ▼ 55 55|Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy ▼ в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard Alignment Number Cells DH DI DJ DK DL DM DN DO DP DR DS DT DU DV 64 65 6 PROJECT W ANSWER a: $442 67 68 69 70 71 72 73 74 75 76 PI= 1 + (NPV/INVESTMENT) W 1.4916 1.4752 Y 15000 Z 1.2409 ANSWER c:Y 8 INVESTMENT IN W & Y TODAY = 900 + 1000-1900 AND NPV = 442.41 + 500 = 942.41, WHICH IS HIGHEST NSWER: PROJECT W &Y 78 79 80 81 82 YEAR 20% IRR(DL80:DR80) 9 IRR 0 4 240 6 240 CF -800 240 240 240 240 NSWER: d : 2096 제 ト 커 | IRRnpv . PBP , external rate | MIRR NpV IRR TS HPR GMAM EAC MACRS LEASE CASH BUDGET BOND EPS EBIT REPLACEMENTWacc bond sensitivil erences: E147 07:16 10-01-2019

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. ▼ ㆆ ▼ Clipboard Alignment Number Cells DH DJ DK DM DN DO DP DR DS DT DU DV 10 ANSWER: d: 83 84 85 86 87 FOR MUTUALLY EXCLUSIVE PROJECTS, NPV IS THE BEST METHOD, IRR NOT RELIABLE YEAR 0 4 6 240 -800 240 240 240 240 240 PAYBACK PERIOD = INVESTMENT/ANNUAL CF-800/240 = 3.33 YEARS 89 YEAR 0 240 1.0000 0.9434 0.8900 0.8396 0.7921 0.7473 0.7050 800.00 226.42 213.60 201.51 190.10 179.34 169.19 800.00 573.58 359.99 158.48 31.63 210.97 380.16 240 240 91 92 93 94 95 96 97 98 CF -800 240 240 240 PV FACTOR AT 6% PV OF CF CUMULATIVE CF DISCOUNTED PAYBACK PERIOD3 YEARS(CCFAT3/CFAT4) IGNORE NEGATIVE SIGN) 3(158.48/190.10) 3.8337 YEARS 100 101 14 ANSWER: DIFFERENCE 3.83-3.33 0.5 YEARS C: 6 MONTHS IRRmpv PBP erences: E147 external rate MIRR NpV IRR REPLACEMENT S i HPR GMAM EAC MACRS LEASE CASH BUDGET i BOND EPS EBIT REPLACEMENT ( vacc bond sens tvt i 07:24 10-01-2019